OpenStudy (anonymous):

3x^+X=10 COMPLETE THE SQUARE

OpenStudy (anonymous):

do you mean this? \[3x^2 + X = 10\]

OpenStudy (softballgirl372015):

i am not sure if this is what you wanted, but i found what x equals.

OpenStudy (anonymous):

I'm not really sure either but I was taught to complete the square by doing \[[\frac{ 1 }{ 2 }b]^2\] With this being said, the a term is 3, the b term is 1, and the c term is 0. So if I did the formula I get this: \[3x^2 + \frac{ 1 }{ 4 }x - 10\frac{ 1 }{ 4 } = 0\]